# Stabilizing Circular Motion Within Nonconcentric Circular Boundary: A Möbius Transformation-Based Approach

### v1

@@ Abstract
Barrier Lyapunov function, circular motion control, Mobius transformation, nonconcentric circular boundary, nonuniform constraints, stabilization.
@@ I Introduction
Dynamical systems, in practice, encounter nonuniform operational constraints. For instance, robotic systems face several such constraints dictated by their application needs, especially in complex environments like irregular, cluttered, or confined spaces, and it is a challenging problem to synthesize a stabilizing controller for their safe operation. Particularly, in geofencing applications where the workspace is specified via virtual boundaries for a real-world geographic area, the robot’s motion is limited by nonuniform spatial constraints, concerning the nature of geofence [1]. Several critical missions such as surveillance or monitoring near territorial boundaries or over a region affected by fire or flood, ocean explorations, autonomous driving in a lane, etc., are some examples where geofencing control of robotic systems is necessary for safety considerations. While allowing operating in the regime of a geofence, it is equally important that the proposed controller also handles nonholonomic motion constraints of a real robotic system such as UAV (unmanned aerial vehicle) and UGV (unmanned ground vehicle).
@@
Motivated by these applications, this paper addresses the problem of stabilizing an underactuated unicycle robot around a desired circular orbit, while restricting its trajectories within a nonconcentric external circular boundary. This differs from conventional circular motion/formation control problems in the literature [2, 3, 4, 5, 6] where no restriction is imposed on the robots’ (or agents’) trajectories. Our previous work [7] in this direction addressed a problem, where the external circular boundary was concentric with the desired stabilizing circular orbit, thereby imposing uniform trajectory-restricting constraints on the robot’s motion due to the constant radial distance between the two circles. An extension to this problem for a general class of polar curves comprising both convex and non-convex curves was studied in [8], where both the stabilizing curve and the external trajectory-constraining boundary were concentric and had a similar nature. As a result, the radial distance between the two curves was constant throughout the complete rotation. However, in practice, the trajectory-constraining set (or boundary) is not necessarily concentric with the desired stabilizing curve, resulting in nonuniform spatial constraints on the robot’s motion. Our consideration of a nonconcentric external circular boundary depicts one such practical scenario and makes the problem challenging.
@@
In recent years, there has been a growing interest among researchers in control design methods that accommodate system constraints, primarily focusing on two solution approaches: control barrier function (CBF)-based methods [9] and barrier Lyapunov function (BLF) based methods [10, 11, 12]. While the former involves solving a quadratic programming (QP) problem for control design, the latter is relatively simpler and extends the concept of conventional control Lyapunov function (CLF) to systems with constraints [13]. Various BLFs have been proposed in the literature to address a range of physical and operational constraints while achieving desired control objectives. These include re-centered barrier function [14], parametric barrier function [15], universal barrier functions [16], tangent-type BLF [17], integral-type BLF [18], among others. In this paper, we adopt the logarithmic-type BLFs, introduced in [10, 11], because of their simplicity in handling the constraints.
@@
The primary challenge in applying BLF-based approaches lies in appropriately designing the so-called barrier functions in their formulation that must accurately characterize the desired constraint requirements of a system or mission. In the literature, two types of barrier functions are commonly discussed: symmetric BLFs and asymmetric BLFs, addressing symmetric and asymmetric system constraints, respectively. These constraints are typically described using (predefined) constant [10] or time-varying barrier functions [11, 16]. When focusing on meeting prescribed transient and steady-state specifications of the system’s response, constraints can be formulated as time-varying barrier functions of output [19]. When aiming to satisfy certain requirements on the system’s states (e.g., the position of a robot in geofencing applications), the barrier functions must be derived considering the constraining-set properties. However, formulating such time-varying barrier functions becomes challenging in arbitrary geofencing regions, where motion constraints exhibit nonuniform spatial dependence and are not solely time-dependent, in general, [20, 21, 22, 23].
@@
To facilitate the subsequent analysis and substantiate various results in this paper, we rely on the following convergence lemma from [10], which is based on the logarithmic BLF.
Lemma 1 (Convergence with BLF[10, pp. 919--920]).
@@ II-A The Möbius Transformation
A Möbius transformation is a mapping of the form [24, Chapter 2, pg. 23], [25, Chapter 3, pg. 137]:
@@
which maps to a constant $a/c$ if $ad-bc=0$. The special case of $c=0$ implies that neither $a$ nor $d$ is zero, and hence, $f(z)$ is well-defined. The quantity $ad-bc$ is called the determinant of $f$. Clearly, the domain of $f$ is $\mathbb{C}\setminus\{-d/c\}$ if $c\neq 0$, and $\mathbb{C}$ if $c=0$. For all $z\in\mathbb{C}\setminus\{-d/c\}$, the derivative of $f$ is given by [24, Chapter 8, pg. 95]:
@@
since $ad-bc\neq 0$. Hence, $f$ is conformal in $\mathbb{C}\setminus\{-d/c\}$, where $c\neq 0$. In the special case of $c=0$, $f(z)=(az+b)/d$ is a linear mapping that is conformal everywhere. In the subsequent discussion, we do not explicitly mention the special case of $c=0$, as it is straightforward. Recall from complex analysis that conformal maps are holomorphic (or analytic). Further, solving for $z$ in (1), the inverse of $f$ is obtained as [24, Chapter 2, pg. 23]:
@@
which has the same determinant $ad-bc\neq 0$ that of $f$, and hence, is also a Möbius transformation. The domain of $f^{-1}$ is $\mathbb{C}\setminus\{a/c\}$. Therefore, $f(z):\mathbb{C}\setminus\{-d/c\}\mapsto\mathbb{C}\setminus\{a/c\}$ defines a bijective and holomorphic map. For simplicity, the Möbius transformation $f$ is often regarded as a mapping from $\mathbb{C}_{\infty}$ to $\mathbb{C}_{\infty}$ such that $f(\infty)=a/c$ and $f(-d/c)=\infty$ if $c\neq 0$, and $f(\infty)=\infty$ if $c=0$, in the extended complex plane $\mathbb{C}_{\infty}=\mathbb{C}\cup\{\infty\}$ (i.e., the complex plane augmented by the point at infinity). Notably, this extension accounts for the “inversion mapping” $z\mapsto 1/z$ on $\mathbb{C}\setminus\{0\}$ to a mapping from $\mathbb{C}_{\infty}$ to $\mathbb{C}_{\infty}$ by the assignments $0\mapsto\infty$ and $\infty\mapsto 0$. With the aid of $\mathbb{C}_{\infty}$, it can be deduced that $f:\mathbb{C}_{\infty}\to\mathbb{C}_{\infty}$ is bijective and conformal everywhere in its domain [24, Chapter 2, pg. 23].
@@ II-B Mapping of Two Nonconcentric Circles to Concentric Circles
The Möbius transformation (1) shares some remarkable properties [25, Chapter 3, pg. 168]: (i) it maps circles to circles (a straight line is viewed as a degenerate circle through infinity), (ii) If two points are symmetric with respect to a circle, then their images under (1) are symmetric with respect to the image circle. This is known as the Symmetry Principle. Note that two points $z$ and $\tilde{z}$ are symmetric with respect to the unit circle if they are related by the inversion mapping $1/\bar{z}$ such that the product of their distances from the center of the circle is unity. This concept can be generalized to a circle with a non-unity radius where this product equals the square of the radius of the circle. For further details, readers are referred to [25, Chapter 3, pp. 139--142]. By leveraging these properties, one can devise a suitable Möbius transformation that maps any two non-intersecting, nonconcentric circles in the $z$-plane to two concentric circles in the $w$-plane. In this direction, we state the following theorem from [26, Chapter 16, pp. 1234--1240] and provide its proof for clarity of further discussion.
@@
Lemma 2 (Non-collinear points [26, Chapter 16, pp. 1234]).
@@ 2. Case II
    Consider that the circle $\mathcal{C}$ encircles circle $\mathcal{C}^{\prime}$ (Fig. 2). In this situation, the hypotheses of Theorem 1 are satisfied only if $\mu<1-|\lambda|$. Since $-1<p<q<1$ in this case, the roots of (13) are again real. Further, these roots are negative (resp., positive) if $p+q>0$ (resp., $p+q<0$), unlike the previous case. Now, following the similar steps above, one can draw the same conclusions in this case as well and are not discussed in detail for brevity.
@@ IV Problem Formulation and Control Design in the transformed plane
In this section, we first restate the problem and design the control in the transformed plane in the spirit of Theorem 1 and Theorem 2 under Möbius transformation (4), which is required to be mapped back to obtain the actual control input $\omega$ in the actual plane. We proceed in this direction as follows:
@@ IV-A System Model and Problem Formulation in Transformed Plane
$\rho=f(r)=\frac{r+\alpha}{r+\beta}=\frac{\alpha(r+\alpha)}{1+\alpha r}\coloneqq|\rho|e^{i\psi},$ (17)
@@
$\dot{\rho}=|\dot{\rho}|e^{i\gamma},\qquad\dot{\gamma}=\Omega,$ (18)
@@
$\mathcal{E}=\rho\pm i|\alpha|e^{i\gamma}=\rho+i\sigma{\rm e}^{i\gamma},$ (19)
where
@@
and the ‘$+$’ and ‘--’ signs are used in the same spirit as in (16). Following Theorem 3, this corresponds to the fact that $\sigma=+|\alpha|$ if $\alpha=\alpha_{s}$, and $\sigma=-|\alpha|$ if $\alpha=\alpha_{\ell}$. The equivalent control objective in the transformed plane is to minimize $\mathcal{E}$. Depending on the root $\alpha_{s}$ or $\alpha_{\ell}$ of (5), Problem 1 can be inferred either as a trajectory-constraining problem or an obstacle-avoidance problem in the transformed plane, respectively. This is illustrated in Fig. 5 where the radial distance $\delta_{T}$ between the two circular boundaries is uniform and is given by
@@
It is important to note here that, analogous to actual error (16) in Fig. 4, the transformed error (19) is defined with respect to mapping $f(\mathcal{C})$ of the desired circular orbit $\mathcal{C}$ for both $\alpha_{s}$ and $\alpha_{\ell}$ (see Fig. 5). The problem can be formally restated in the transformed plane as:
@@ IV-B Control Design in Transformed Plane
˙$\dot{\mathcal{E}}=\dot{\rho}-\sigma{\rm e}^{i\gamma}\dot{\gamma}=|\dot{\rho}|e^{i\gamma}-\sigma{\rm e}^{i\gamma}\dot{\gamma}=(|\dot{\rho}|-\sigma\Omega){\rm e}^{i\gamma},$ (24)
@@
$\dot{\mathcal{S}}=\frac{\langle\rho,e^{i\gamma}\rangle}{\delta_{T}^{2}-|\mathcal{E}|^{2}}(|\dot{\rho}|-\sigma\Omega).$ (25)
@@
    $\dot{\mathcal{S}}=-\kappa\left[\frac{\langle\rho,e^{i\gamma}\rangle}{\delta_{T}^{2}-|\mathcal{E}|^{2}}\right]^{2}\leq 0,$ (28)
@@
    $\frac{\langle\rho,e^{i\gamma}\rangle}{\delta_{T}^{2}-|\mathcal{E}|^{2}}=0\implies\langle\rho,e^{i\gamma}\rangle=0,$ (29)
    in $\Gamma$. Substituting (29) into (26) implies that $\Omega=|\dot{\rho}|/\sigma$ in the set $\Gamma$. Consequently, it follows from (24) that ˙$\dot{\mathcal{E}}=0$ in the set $\Gamma$. Now, substituting for $\rho$ from (17) into (29), results in $\langle\mathcal{E}-i\sigma{\rm e}^{i\gamma},{\rm e}^{i\gamma}\rangle=0\implies\langle\mathcal{E},{\rm e}^{i\gamma}\rangle-\sigma\langle i{\rm e}^{i\gamma},e^{i\gamma}\rangle=0\implies\langle\mathcal{E},{\rm e}^{i\gamma}\rangle=0$, since $\langle i{\rm e}^{i\gamma},{\rm e}^{i\gamma}\rangle=0$. For $\langle\mathcal{E},{\rm e}^{i\gamma}\rangle=0$ to hold, it is needed that $\mathcal{E}=0$, as ˙$\dot{\mathcal{E}}=0$ in $\Gamma$. This, together with the fact that $\Omega=|\dot{\rho}|/\sigma$ in $\Gamma$, implies that the trajectories of (18) asymptotically converge around the circle of radius $|\sigma|=|\alpha|$ with the linear speed $|\dot{\rho}|$ and the direction of rotation determined by $\sigma$ (see Theorem 3). Alternatively, $\mathcal{E}\to 0$ in the set $\Gamma\subset\Xi\subset\mathcal{W}_{\delta_{T}}$, as $t\to\infty$.

### v2

# Möbius Transformation-Based Circular Motion Control for Unicycle Robots in Nonconcentric Circular Geofences

@@ Abstract
Möbius transformation, geofencing control, nonuniform constraints, circular motion control, unicycle robot, barrier Lyapunov function, stabilization.
@@ I Introduction
Dynamical systems, in practice, encounter nonuniform operational constraints. For instance, robotic systems face several such constraints dictated by their application needs, especially in complex environments like irregular, cluttered, or confined spaces, and it is a challenging problem to synthesize a stabilizing controller for their safe operation. Particularly, in geofencing applications where the workspace is specified via virtual boundaries for a real-world geographic area, the robot’s motion is limited by nonuniform spatial constraints, concerning the nature of geofence [1]. Several critical missions such as surveillance or monitoring near territorial boundaries or over a region affected by fire or flood, ocean explorations, autonomous driving in a lane, etc., are some examples where geofencing control of robotic systems is necessary for safety considerations. While allowing operating in the regime of a geofence, it is equally important that the proposed controller also handles nonholonomic motion constraints of a real robotic system such as UAV (unmanned aerial vehicle) and UGV (unmanned ground vehicle) [2].
@@
Motivated by these applications, this paper addresses the problem of stabilizing an underactuated unicycle robot around a desired circular orbit while restricting its trajectories within a nonconcentric external circular boundary. This differs from conventional circular motion/formation control problems in the literature [3, 4, 5, 6, 7, 8] where no restriction is imposed on the robots’ (or agents’) trajectories. Our previous work [9] in this direction addressed a problem, where the external circular boundary was concentric with the desired stabilizing circular orbit, thereby imposing uniform trajectory-restricting constraints on the robot’s motion due to the constant radial distance between the two circles. An extension to this problem for a general class of polar curves was studied in [10], where both the stabilizing curve and the external trajectory-constraining boundary were concentric and had a similar nature. As a result, the radial distance between the two curves was constant throughout the complete rotation. However, in practice, the trajectory-constraining set (or boundary) is not necessarily concentric with the desired stabilizing curve, resulting in nonuniform spatial constraints on the robot’s motion. Our consideration of a nonconcentric external circular boundary depicts one such practical scenario and makes the problem challenging.
@@
In recent years, there has been a growing interest among researchers in control design methodologies that accommodate system and operational constraints, primarily focusing on two solution approaches: control barrier function (CBF)-based methods [11, 12] and barrier Lyapunov function (BLF) based methods [13, 14]. While the former involves solving a quadratic programming (QP) problem for control design [15], the latter is relatively simpler and extends the concept of conventional control Lyapunov function (CLF) to systems with constraints [16]. Various BLFs have been proposed in the literature to address a range of physical and operational constraints while achieving desired control objectives. These include re-centered barrier function [17], parametric barrier function [18], universal barrier functions [19], tangent-type BLF [20], integral-type BLF [21], among others. In this paper, we adopt the logarithmic-type BLF, introduced in [13], because of its simplicity in handling the constraints.
@@
The primary challenge in applying BLF-based approaches lies in appropriately designing the so-called barrier functions in their formulation that must accurately characterize the desired constraint requirements of a system or mission. In the literature, two types of barrier functions are commonly discussed: symmetric BLFs and asymmetric BLFs, addressing symmetric and asymmetric system constraints, respectively [13, 19]. When focusing on meeting prescribed transient and steady-state specifications of the system’s response, constraints can be formulated as time-varying barrier functions of output [22]. When aiming to satisfy certain requirements on the system’s states (e.g., the position of a robot in geofencing applications), the barrier functions must be derived considering the constraining-set properties. However, formulating such time-varying barrier functions becomes challenging in arbitrary geofencing regions, where motion constraints exhibit nonuniform spatial dependence and are not solely time-dependent, in general, [23, 24, 25, 26].
@@
To facilitate the subsequent analysis and substantiate various results in this paper, we state the following definition and a convergence lemma from [13], characterizing the properties of BLF.
Definition 1 (Barrier Lyapunov Function [13]).
A Barrier Lyapunov Function is a scalar function $V(\mathcal{X})$ of state vector $\mathcal{X}\in\mathcal{D}$ of the system $\dot{\mathcal{X}}=\mathfrak{f}(\mathcal{X})$ on an open region $\mathcal{D}$ containing the origin, that is continuous, positive-definite, has continuous first-order partial derivatives at every point of $\mathcal{D}$, has the property $V(\mathcal{X})\to\infty$ as $\mathcal{X}$ approaches the boundary of $\mathcal{D}$, and satisfies $V(\mathcal{X}(t))\leq\varphi_{0},\forall t\geq 0$, along the solution of $\dot{\mathcal{X}}=\mathfrak{f}(\mathcal{X})$ for $\mathcal{X}(0)\in\mathcal{D}$ and some positive constant $\varphi_{0}$.
Lemma 1 (Convergence with BLF[13, pp. 919--920]).
@@ II-A The Möbius Transformation
A Möbius transformation is a mapping of the form [27, Chapter 2, pg. 23], [28, Chapter 3, pg. 137]:
@@
which maps to a constant $a/c$ if $ad-bc=0$. The special case of $c=0$ implies that neither $a$ nor $d$ is zero, and hence, $f(z)$ is well-defined. The quantity $ad-bc$ is called the determinant of $f$. Clearly, the domain of $f$ is $\mathbb{C}\setminus\{-d/c\}$ if $c\neq 0$, and $\mathbb{C}$ if $c=0$. For all $z\in\mathbb{C}\setminus\{-d/c\}$, the derivative of $f$ is given by [27, Chapter 8, pg. 95]:
@@
since $ad-bc\neq 0$. Hence, $f$ is conformal in $\mathbb{C}\setminus\{-d/c\}$, where $c\neq 0$. In the special case of $c=0$, $f(z)=(az+b)/d$ is a linear mapping that is conformal everywhere. In the subsequent discussion, we do not explicitly mention the special case of $c=0$, as it is straightforward. Recall from complex analysis that conformal maps are holomorphic (or analytic). Further, solving for $z$ in (1), the inverse of $f$ is obtained as [27, Chapter 2, pg. 23]:
@@
which has the same determinant $ad-bc\neq 0$ that of $f$, and hence, is also a Möbius transformation. The domain of $f^{-1}$ is $\mathbb{C}\setminus\{a/c\}$. Therefore, $f(z):\mathbb{C}\setminus\{-d/c\}\mapsto\mathbb{C}\setminus\{a/c\}$ defines a bijective and holomorphic map. For simplicity, the Möbius transformation $f$ is often regarded as a mapping from $\mathbb{C}_{\infty}$ to $\mathbb{C}_{\infty}$ such that $f(\infty)=a/c$ and $f(-d/c)=\infty$ if $c\neq 0$, and $f(\infty)=\infty$ if $c=0$, in the extended complex plane $\mathbb{C}_{\infty}=\mathbb{C}\cup\{\infty\}$ (i.e., the complex plane augmented by the point at infinity). Notably, this extension accounts for the “inversion mapping” $z\mapsto 1/z$ on $\mathbb{C}\setminus\{0\}$ to a mapping from $\mathbb{C}_{\infty}$ to $\mathbb{C}_{\infty}$ by the assignments $0\mapsto\infty$ and $\infty\mapsto 0$. With the aid of $\mathbb{C}_{\infty}$, it can be deduced that $f:\mathbb{C}_{\infty}\to\mathbb{C}_{\infty}$ is bijective and conformal everywhere in its domain [27, Chapter 2, pg. 23].
@@ II-B Mapping of Two Nonconcentric Circles to Concentric Circles
The Möbius transformation (1) shares some remarkable properties [28, Chapter 3, pg. 168]: (i) it maps circles to circles (a straight line is viewed as a degenerate circle through infinity), (ii) If two points are symmetric with respect to a circle, then their images under (1) are symmetric with respect to the image circle. This is known as the symmetry principle. Note that two points $z$ and $\tilde{z}$ are symmetric with respect to the unit circle if they are related by the inversion mapping $1/\bar{z}$ such that the product of their distances from the center of the circle is unity. This concept can be generalized to a circle with a non-unity radius where this product equals the square of the radius of the circle. For further details, readers are referred to [28, Chapter 3, pp. 139--142]. By leveraging these properties, one can devise a suitable Möbius transformation that maps any two non-intersecting, nonconcentric circles in the $z$-plane to two concentric circles in the $w$-plane. In this direction, we state the following theorem from [29, Chapter 16, pp. 1234--1240] and provide its proof for clarity of further discussion.
@@
Lemma 2 (Non-collinear points [29, Chapter 16, pp. 1234]).
@@ 2. Case II
    Consider that the circle $\mathcal{C}$ encircles circle $\mathcal{C}^{\prime}$ (Fig. 2). In this situation, the hypotheses of Theorem 1 are satisfied only if $\mu<1-|\lambda|$. Since $-1<p<q<1$ in this case, the roots of (13) are again real. Further, these roots are negative (resp., positive) if $p+q>0$ (resp., $p+q<0$), unlike the previous case. Now, following the similar steps as above, one can draw the same conclusions in this case as well and are not discussed in detail for brevity.
@@ IV Problem Formulation and Control Design in the transformed plane
In this section, we begin by restating the problem and designing the control in the transformed plane, following Theorem 1 and Theorem 2 under the Möbius transformation (4). This control design needs to be mapped back to the actual plane to obtain the real control input $\omega$. We proceed in this direction as follows:
@@ IV-A System Model and Problem Formulation in Transformed Plane
$\rho=f(r)=\frac{r+\alpha}{r+\beta}=\frac{\alpha(r+\alpha)}{1+\alpha r}\coloneqq|\rho|{\rm e}^{i\psi},$ (17)
@@
$\dot{\rho}=|\dot{\rho}|{\rm e}^{i\gamma},\qquad\dot{\gamma}=\Omega,$ (18)
@@
$\mathcal{E}=\rho\pm i|\alpha|{\rm e}^{i\gamma}=\rho+i\sigma{\rm e}^{i\gamma},$ (19)
where,
@@
and the ‘$+$’ and ‘--’ signs are used in the same spirit as in (16). Following Theorem 3, this corresponds to the fact that $\sigma=+|\alpha|$ if $\alpha=\alpha_{s}$, and $\sigma=-|\alpha|$ if $\alpha=\alpha_{\ell}$. The equivalent control objective in the transformed plane is to minimize $\mathcal{E}$ in (19). Depending on the root $\alpha_{s}$ or $\alpha_{\ell}$ of (5), Problem 1 can be inferred either as a trajectory-constraining problem or an obstacle-avoidance problem in the transformed plane, respectively. This is illustrated in Fig. 5 where the radial distance $\delta_{T}$ between the two circular boundaries is uniform and is given by
@@
It is important to note that, analogous to actual error (16) in Fig. 4, the transformed error (19) is defined with respect to mapping $f(\mathcal{C})$ of the desired circular orbit $\mathcal{C}$ for both $\alpha_{s}$ and $\alpha_{\ell}$ (see Fig. 5). The problem can be formally restated in the transformed plane as:
@@ IV-B Control Design in Transformed Plane
˙$\dot{\mathcal{E}}=\dot{\rho}-\sigma{\rm e}^{i\gamma}\dot{\gamma}=|\dot{\rho}|{\rm e}^{i\gamma}-\sigma{\rm e}^{i\gamma}\dot{\gamma}=(|\dot{\rho}|-\sigma\Omega){\rm e}^{i\gamma},$ (24)
@@
$\dot{\mathcal{S}}=\frac{\langle\rho,{\rm e}^{i\gamma}\rangle}{\delta_{T}^{2}-|\mathcal{E}|^{2}}(|\dot{\rho}|-\sigma\Omega).$ (25)
@@
    $\dot{\mathcal{S}}=-\kappa\left[\frac{\langle\rho,{\rm e}^{i\gamma}\rangle}{\delta_{T}^{2}-|\mathcal{E}|^{2}}\right]^{2}\leq 0,$ (28)
@@
    $\frac{\langle\rho,{\rm e}^{i\gamma}\rangle}{\delta_{T}^{2}-|\mathcal{E}|^{2}}=0\implies\langle\rho,{\rm e}^{i\gamma}\rangle=0,$ (29)
    in $\Gamma$. Substituting (29) into (26) implies that $\Omega=|\dot{\rho}|/\sigma$ in the set $\Gamma$. Consequently, it follows from (24) that ˙$\dot{\mathcal{E}}=0$ in the set $\Gamma$. Now, substituting for $\rho$ from (17) into (29), results in $\langle\mathcal{E}-i\sigma{\rm e}^{i\gamma},{\rm e}^{i\gamma}\rangle=0\implies\langle\mathcal{E},{\rm e}^{i\gamma}\rangle-\sigma\langle i{\rm e}^{i\gamma},{\rm e}^{i\gamma}\rangle=0\implies\langle\mathcal{E},{\rm e}^{i\gamma}\rangle=0$, since $\langle i{\rm e}^{i\gamma},{\rm e}^{i\gamma}\rangle=0$. For $\langle\mathcal{E},{\rm e}^{i\gamma}\rangle=0$ to hold, it is needed that $\mathcal{E}=0$, as ˙$\dot{\mathcal{E}}=0$ in $\Gamma$. This, together with the fact that $\Omega=|\dot{\rho}|/\sigma$ in $\Gamma$, implies that the trajectories of (18) asymptotically converge around the circle of radius $|\sigma|=|\alpha|$ with the linear speed $|\dot{\rho}|$ and the direction of rotation determined by $\sigma$ (see Theorem 3). Alternatively, $\mathcal{E}\to 0$ in the set $\Gamma\subset\Xi\subset\mathcal{W}_{\delta_{T}}$, as $t\to\infty$.